OpenStudy (anonymous):

I have the answer can somone please check me. What is the derivative of m(x)=ktan(kx) where k is the constant

OpenStudy (anonymous):

answer will be\[k^2(1+\tan^2 kx)\]is that what u got?

OpenStudy (anonymous):

I got k^2cos^2(kx) is thatthe same thing

OpenStudy (anonymous):

emm we have a formula as\[1+\tan^2 x=\frac{1}{\cos^2 x}\]so u will come up with\[\frac{k^2}{\cos^2 kx}\]

OpenStudy (anonymous):

are u familiar with that formula? :)

OpenStudy (anonymous):

yes ok I gotcha no thanks

OpenStudy (anonymous):

ok :)